Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Sat Nov 28 20:43:19 MET 1998) Success
No known problems

Onsala (John Conway at Sun Nov 29 20:45:29 MET 1998) Minor failures
No known problems

Torun (K.Borkowski at Mon Nov 30 10:30:06 MET 1998) Success
No problems reported.

Noto (C.Trigilio at Mon Nov 30 11:16:57 MET 1998) Success

Effelsberg (dgraham at Mon Nov 30 15:16:18 MET 1998) Minor failures
OK. Calibration will need care , as Tcal obviously different LCP/RCP

Jodrell (Alastair Gunn at Tue Dec 1 08:12:34 MET 1998) Minor failures
GPS clock appeared to stop at 15.31, restarted on time at 15.44. Some problems because PI did not provide adequate slew times for the Lovell telescope. 1613+341 scan at 16:00 missed, J0958+65 scan at 16:22 missed, on M82 at 16:28.

Cambridge (Alastair Gunn at Tue Dec 1 08:13:11 MET 1998) Failure
Experiment cancelled for Cambridge 32-m telescope due to serious hardware fault with VLBA recorder head stack.

Westerbork (Tony Foley at Tue Dec 1 18:30:32 MET 1998) Minor failures
Used single-dish setup instead of tied-array. Tsys in IF1 about 28K and IF2 about 40K with 0.096K/Jy. Tsys determinations made at the time were very noisy, but it is not clear why.

Calibration (Jean\-Francois Desmurs at Tue Dec 15 08:59:48 MET 1998) Success
(Jb unstable values on VC03 and 04) The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gp018.antab. Read the comments in the file to get further information.